OpenStudy (anonymous):

x^2-7x= - 24/4 please solve for x

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

use the quadratic equation to find the roots. once you get used to it, you can find that the roots are x=1 or x=6

OpenStudy (anonymous):

and the quadratic equation would be?

OpenStudy (anonymous):

ok, so i said that a little funky. x^2-7x=-24/4 is a quadratic equation. you can rearrange it to a simpler form of x^2-7x+6=0. From that, you can find that that can be rearranged to the equation (x-6)(x-1)=0. The roots are then easy to see.

OpenStudy (anonymous):

thank you:) that second explanation helped alot.. im a bit slow sorry

OpenStudy (anonymous):

no problem. alternatively, you can use the equation \[ x= -b \pm (\sqrt{b^2-4ac})/(2a)\] where a=1, b=-7 and c=6. but that takes a lot longer and is kind of a last resort if you can't see how to rearrange it right away.

OpenStudy (anonymous):

well this next question might be a bit irrelevant to the above questions but: x^2 - 7x what number would i add to the polynomial to complete the sqaure? the answer i got was 12.25 is that correct?

OpenStudy (anonymous):

that seems right to me.